### 02 Apr 2026 v1

# RFOX (Rotated-Field Oscillatory eXchange) quantum algorithm: Towards Parameter-Free Quantum Optimizers

Abstract
We introduce RFOX (Rotated-Field Oscillatory eXchange), a parameter-free quantum algorithm for combinatorial optimization. RFOX combines an almost constant non-stoquastic $XX$ catalyst with a weak harmonic $ZX$ counter-diabatic term. Using the Floquet-Magnus expansion, we derive a closed-form effective Hamiltonian whose first-order term retains the full $XX$ driver, while the leading correction consists of a single qubit $Y$ field at high drive frequency. This structure ensures that the instantaneous spectral gap remains essentially flat, independent of both the interpolation parameter and the disorder strength, modulated only by a $\delta$ parameter. This behavior stands in stark contrast to the unpredictable gap reductions, or even collapses, exhibited by the XX (stoquastic), $XX$, and $X+sXX$ (non-stoquastic) driver schedules. Extensive noiseless simulations on random-field Ising model (RFIM) instances with 7, 9, and 12 qubits, across three magnetic-field ranges, validate these spectral predictions: RFOX attains near-optimal, and in some cases exact, ground states using up to an order of magnitude fewer Trotter slices. Its performance advantage grows with increasing disorder, as conventional methods slow down near vanishing gaps, whereas RFOX maintains a constant runtime scaling of $T\propto\Delta_{\min}^{-2}$. Hardware experiments on IBM Quantum processors (Eagle r3 and Heron r1, with 12, 15, and 20 physical qubits) reproduce similar performance rankings. These results suggest that fixed-gap, non-stoquastic drivers augmented with analytically derived counter-diabatic terms offer a promising, scalable, and tuning-free route toward quantum optimizers for combinatorial optimization problems.
I Introduction
Inside the quantum computing paradigm, the Variational Quantum Algorithms (VQAs) have become a cornerstone approach for harnessing near-term quantum devices in the noisy intermediate-scale quantum (NISQ) era. Landmark schemes such as the Variational Quantum Eigensolver (VQE) Peruzzo et al. (2014b); Fedorov et al. (2022); Liu et al. (2022) and the Quantum Approximate Optimization Algorithm (QAOA) Farhi et al. (2014); Blekos et al. (2024) exemplify how shallow, parameterized circuits, optimized through a quantum-classical feedback loop, can already tackle demanding problems in quantum chemistry and combinatorial optimization that strain classical resources Zhou et al. (2020, 2023); Crooks (2018); Bae and Lee (2024). Beyond these initial demonstrations, a rapidly expanding body of work continues to expand the reach of VQAs, driving progress in ansatz engineering, error-resilient circuit design, and strategies for effective parameter initialization Preskill (2018); Peruzzo et al. (2014a); Cerezo et al. (2021); Motta et al. (2020).
Despite these successes, the hybrid nature of VQAs imposes fundamental bottlenecks. As the system size increases, the classical optimizer encounters barren plateaus, regions where cost-function gradients vanish exponentially, which delay training even in noiseless simulations Grover (1996); Schuld et al. (2014); Haug and Kim (2021); Zhang et al. (2022). In real hardware, stochastic gate errors further flatten the optimization landscape and introduce shot noise, overwhelming gradient signals, and making large-scale optimization prohibitively expensive in both circuit executions and classical processing time McClean et al. (2018); Wang et al. (2021).
To address these limitations, we introduce Rotated-Field Oscillatory eXchange (RFOX) quantum algorithm, an entirely quantum protocol tailored to the random-field Ising model (RFIM) problem. Instead of alternating between quantum circuits and a classical optimizer, RFOX directly encodes the local fields via phase rotations and unifies the cost and mixing dynamics within a single two‐qubit gate sequence, $U(\theta,\phi)=RZX(\theta)\ RXX(\phi)$, applied to each edge of the RFIM graph. By setting $\theta,\phi\propto\Delta t$, this construction implements a first‐order Trotter step inspired by the quantum adiabatic theorem. To enhance adiabaticity, both gates are modulated by a small harmonic amplitude $\delta\ll 1$, synchronized with $\Delta t$, generating counter‐diabatic corrections while preserving an almost constant non-stoquastic gap.
In disordered systems such as the RFIM, where frustrated couplings and random longitudinal fields hinder conventional quantum annealing, introducing a transverse interaction channel $X\otimes X$ (XX) can accelerate convergence and suppress localization-induced gap closures, outperforming pure transverse field schedules Suzuki et al. (2007); Zhou and others (2022); Yang and others (2024). A key design element of RFOX is the combination of a nearly constant non‐stoquastic XX driver with a counter-diabatic term of time oscillation $Z\otimes X$ (ZX). The XX interaction alleviates the stoquastic bottleneck by coupling computational basis states differing by multiple spin flips, thus widening the spectral gap Lykiardopoulou et al. (2021); Albash (2019); Choi (2021). However, a large gap alone does not guarantee adiabatic evolution: when problem fields vary rapidly, residual diabatic transitions can accumulate. The harmonic ZX term serves as the leading-order component of the adiabatic gauge potential, modulated by a phase-shifted sinusoidal envelope that supplies precisely the time derivative required to cancel non-adiabatic couplings, without adding any static cost term that could close the gap. As a result, the combined XX+ZX driver preserves a nearly constant minimum gap while suppressing diabatic leakage, enabling the system to reach approximate, and in some cases exact, ground states at fixed, shallow depth and without classically tuned parameters.
Simulations in 7-12 spin RFIM instances show that a circuit depth of $p=100$ consistently yields near-optimal, and often exact, ground states, outperforming all other tested methods. We further examine the convergence and scalability of RFOX in randomly generated RFIM instances with 12 to 20 qubits executed on IBM Quantum backends: ibm_brisbane, ibm_sherbrooke, and ibm_torino. Compared with conventional drivers, either an XX-only catalyst or an X + sXX schedule, RFOX achieves higher ground-state fidelities and faster convergence. By eliminating the classical optimization loop, the protocol reduces operational overhead, improving both the practicality and scalability of quantum optimization on near-term hardware.
Section II reviews the theoretical background of the algorithm. Section III provides a formal description of RFOX. Section IV details the generation of RFIM instances and the implementation of these problems on quantum hardware. Section V presents our simulation and experimental results, and Section VI concludes the paper.
II Theoretical Background
In this section, we outline the main concepts underlying the RFOX quantum algorithm, which will be detailed in the following section.
RFIM problems can be represented as an undirected graph $G=(V,E)$ with $|V|=N$ sites (or as a $d$-dimensional lattice when $G$ is regular), where each classical Ising spin $\sigma_{i}\in\{-1,+1\}$ for $i\in V$. The RFIM Hamiltonian Cipra (1987); Hughes (1999) is:
$H=-\sum_{\langle i,j\rangle\in E}J_{ij}\sigma_{i}\sigma_{j}-\sum_{i\in V}h_{i}\sigma_{i}$ (1)
Here, $J_{ij}$ denotes the coupling strength, positive for ferromagnetic interactions ($J_{ij}>0$) and negative for antiferromagnetic interactions ($J_{ij}<0$), $h_{i}$ are independent magnetic fields, and $\langle i,j\rangle$ indicates nearest-neighbor pairs.
This Hamiltonian has two energy contributions:
  • The interaction energy $\sum_{\langle i,j\rangle\in E}J_{ij}\sigma_{i}\sigma_{j}$, favors aligning spins to minimize the pairwise coupling energy.
  • The field energy $\sum_{i\in V}h_{i}\sigma_{i}$, favors each spin aligning with its local random field.
II.1 Magnetic field encoding using phase gates
In RFOX, we develop a phase mapping of the magnetic field energy. The term $\sum_{j\in V}h_{j}\,\sigma_{j}$ is mapped to phase rotations $P(\phi_{j})$ via a sequence of phase gates that create an interference pattern between two Walsh-Hadamard transforms:
$H_{B}=H^{\otimes n}P(\phi_{j})H^{\otimes n}=e^{-i\frac{\phi_{j}}{2}HZH}\ .$ (2)
Here, $H^{\otimes n}$ denotes the Walsh-Hadamard transform over $n$ qubits. See Appendix A for details.
This phase encoding generates an initial candidate ground state (or states) by quantum interference, accounting only for the field energy contribution.
II.2 Stoquastic vs. non-stoquastic Hamiltonians
To encode the interaction energy term from RFIM in RFOX, we leverage principles from the adiabatic theorem, quantum walks, nonstoquastic Hamiltonians, and counter-diabatic driving.
A stoquastic Hamiltonian is one that can be written (on some basis) such that all elements of the off-diagonal matrix are real and non-positive, whereas a non-stoquastic Hamiltonian violates this condition by allowing some (or all) off-diagonal entries to be positive or non-real in any choice of local basis. This sign structure has some important consequences:
  • Quantum Monte Carlo simulations are free from the negative sign problem, making stoquastic models tractable on classical hardware Ceperley and Alder (1986); Austin et al. (2012); Gubernatis et al. (2016).
  • In many-body adiabatic algorithms, the Perron–Frobenius theorem guarantees strictly positive ground-state amplitudes for stoquastic Hamiltonians, which suppresses destructive interference and can hinder tunneling through tall, narrow energy barriers Bravyi et al. (2006); Bravyi (2014); Klassen and Terhal (2019).
  • Non-stoquastic Hamiltonians may exhibit ground-state amplitudes with mixed signs, enabling destructive interference paths and enhanced tunneling through rugged cost landscapes Hormozi et al. (2017a).
RFOX exploits a non-stoquastic XX interaction Marvian et al. (2018); Klassen et al. (2020); Ioannou et al. (2020) to allow ground-state amplitudes with mixed signs, thereby enabling stronger quantum interference. In practice, such non-stoquastic drivers can enlarge the minimum adiabatic gap and improve the optimization performance Vinci and Lidar (2017); Nishimori and Takada (2017). In fact, non-stoquastic XX drivers have been shown to enhance success probabilities on challenging spin-glass and RFIM instances, sometimes exponentially, relative to standard transverse field protocols Hormozi et al. (2017b); Lykiardopoulou et al. (2021); Albash (2019); Choi (2021).
After encoding the magnetic field phase, we model the RFIM interaction using both $RXX(\beta)$ and $RZX(\gamma)$ operators. Rather than a continuous, time-dependent evolution, we divide the circuit into $p$ discrete slices $k=0,\dots,p-1$. We define:
$H_{XX}(k)=A_{XX}(k)\sum_{(u,v)\in E}X_{u}X_{v},\quad A_{XX}(k)=1-\delta\cos\bigl(2\pi Nk/p\bigr),$ (3)
where $E$ is the edge set and $N$ is the number of qubits. Key features of $H_{XX}(k)$ include:
  • In the computational basis, $X_{u}X_{v}$ flips qubits $u$ and $v$ simultaneously; its off-diagonal entries take both positive and negative values, making $H_{XX}$ non-stoquastic.
  • The almost constant part ($A_{XX}\approx 1$) keeps the driver always active, while the small harmonic modulation $\delta\ll 1$ synchronizes with the high-frequency ZX kicks, providing counter-diabatic assistance.
For the ZX term, we define its action as:
$H_{ZX}(k)=B_{ZX}(k)\sum_{(u,v)\in E}Z_{u}X_{v},\quad B_{ZX}(k)=\delta\sin\bigl(2\pi Nk/p\bigr).$ (4)
Here, $Z_{u}X_{v}$ flips qubit $v$ and multiplies its amplitude by $\pm 1$ depending on the state of qubit $u$. This operator has three key effects:
  • Since $B_{ZX}(k)$ is antisymmetric with zero mean, it serves as a first-order counter-diabatic (CD) correction, canceling diabatic transitions induced by $H_{XX}$ to leading order without altering the instantaneous spectrum Cárdenas-López and Chen (2022); Claeys et al. (2019).
  • Its mixed-sign off-diagonal elements lift near-degeneracies, widening the minimum energy gap and improving success probabilities at finite circuit depth Chandarana et al. (2024).
  • The factor of $Z_{u}$ ties each kick’s amplitude to the current spin configuration, biasing collective spin flips toward low‑energy valleys in the RFIM landscape.
Because $\sum_{k=0}^{p-1}B_{ZX}(k)=0$, the total number of slices remains $p$, and the circuit depth scales as $2p\lvert E\rvert$, independent of any adiabatic‐schedule parameter.
II.3 Counter‑diabatic driving and the adiabatic gauge potential
Counter-diabatic (CD), or transitionless, driving augments a time-dependent Hamiltonian with an auxiliary term that suppresses non-adiabatic transitions, enabling the system to remain in its instantaneous eigenstate throughout the evolution Cárdenas-López and Chen (2022); Claeys et al. (2019). This term is formally given by the adiabatic gauge potential (AGP), which encodes the geometric curvature of the eigenstates as the control parameter $\lambda$ varies Claeys et al. (2019).
For a Hamiltonian $H(\lambda)$ with control parameter $\lambda(t)$, the AGP $A_{\lambda}$ satisfies (in units $\hbar=1$)
$A_{\lambda},H(\lambda)]=i\partial_{\lambda}H(\lambda)\ .$ (5)
In the instantaneous eigenbasis $\{|n(\lambda)\rangle\}$, $A_{\lambda}$ can be written as
$A_{\lambda}=i\sum_{m\neq n}\frac{\langle m|\partial_{\lambda}H|n\rangle}{E_{n}-E_{m}}|m\rangle\langle n|\ ,$ (6)
producing off-diagonal couplings that exactly cancel the diabatic excitations. For spin systems, $A_{\lambda}$ is an imaginary Hermitian operator and typically contains an odd number of Pauli‑$Y$ matrices Sels and Polkovnikov (2017).
RFOX implements this CD mechanism via its time‑dependent ZX operator. Specifically:
  • The second-order Magnus term ($i\,\frac{\delta^{2}}{\omega}\sum_{u}Y_{u}$) emerges naturally from the fast, harmonic ZX drive and matches the local AGP (AGP consistency) for an Ising-type system driven by a transverse field. The single $Y$ Pauli terms anticommutes with both the transverse-field term in $H_{B}$ and the XX backbone, providing the minimal transition set needed to cancel diabatic excitations.
  • Instead of adding $\sum_{u}Y_{u}$ explicitly, RFOX synthesises the CD term: the $\sin(\omega t)ZX$ kick, combined with the static XX interaction, generates $i\frac{\delta^{2}}{\omega}\sum_{u}Y_{u}$ via the commutator $[XX,ZX]$ in the Magnus expansion. This produces the CD correction without enlarging the native control set.
  • The Floquet‑engineered $Y$ term prevents gap closures along the evolution path by converting would-be level crossings into avoided crossings. Numerical simulations and prior CD-engineering studies confirm that this mechanism maintains a near-constant minimum gap and high ground-state fidelity at finite runtimes Cárdenas-López and Chen (2022); Claeys et al. (2019); Sels and Polkovnikov (2017).
III RFOX quantum algorithm description
In this section, we present the theory behind the RFOX algorithm. We begin with the magnetic-field encoding using phase gates, then describe how two-qubit interactions are modeled via XX and ZX gates modulated by harmonic functions to evolve the system toward its ground-state.
III.1 Magnetic-field energy phase encoding
The magnetic-field encoding is implemented by generating an interference pattern. We begin by preparing an equal superposition of all computational basis states via a sequence of Hadamard gates,
$|\psi\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{x=0}^{2^{n}-1}|x\rangle.$ (7)
Each magnetic-field value $h_{i}$ is then mapped to a corresponding phase according to
$\phi_{i}=\pi\ \frac{h_{i}/|h|_{\max}+1}{2}.$ (8)
Under this mapping, negative magnetic-field values correspond to phases closer to 0, whereas positive values correspond to phases closer to $\pi$. This transformation is non-trivial, as it embeds the magnetic energy of the problem directly into the quantum state’s phase structure. After the phase encoding, a second sequence of Hadamard gates transforms the phase information back to the $Z$ basis, enhancing the probability amplitudes of states whose spin configurations align more favorably with the applied magnetic field (see Appendix A for a detailed derivation).
The corresponding Hamiltonian for the magnetic-field phase encoding is expressed as
$H_{B}=\sum_{j=0}^{N-1}\phi_{j}X_{j}\ ,$ (9)
where $\phi_{j}=\pi\ (h_{j}/|h|_{\max}+1)/2$ encodes the magnetic-field values.
This encoding scheme is particularly suited for the RFOX framework because it directly incorporates the problem’s magnetic-field energy landscape into the initial quantum state. Unlike the uniform phase-separation steps in QAOA, this method scales the applied phases according to the relative intensity of the local field, biasing the initial state toward energetically favorable spin configurations. By steering the evolution toward low-energy regions of the configuration space from the outset, the algorithm achieves faster convergence, reduced dependence on classical post-processing, and improved robustness in NISQ devices where circuit depth is limited.
III.2 Interaction energy modeling
To model the interaction between two spins in random field Ising model, RFOX uses a combination of operators.
$RXX(\theta)=e^{-i\frac{\theta}{2}X\otimes X}=cos\left(\frac{\theta}{2}\right)\mathbb{I}_{4}-i\ sin\left(\frac{\theta}{2}\right)X\otimes X\ ,$ (10)
and
$RZX(\phi)=e^{-i\frac{\phi}{2}Z\otimes X}=cos\left(\frac{\phi}{2}\right)\mathbb{I}_{4}-i\ sin\left(\frac{\phi}{2}\right)Z\otimes X\ ,$ (11)
gates. When analyzing the resulting matrix operator $U(\theta,\phi)=RZX(\theta)RXX(\phi)$:
$U(\theta,\phi)=R_{ZX}(\theta)\,R_{XX}(\phi)=\begin{pmatrix}c_{\theta}&-is_{\theta}&0&0\\ -is_{\theta}&c_{\theta}&0&0\\ 0&0&c_{\theta}&is_{\theta}\\ 0&0&is_{\theta}&c_{\theta}\end{pmatrix}\begin{pmatrix}c_{\phi}&0&0&-is_{\phi}\\ 0&c_{\phi}&-is_{\phi}&0\\ 0&-is_{\phi}&c_{\phi}&0\\ -is_{\phi}&0&0&c_{\phi}\end{pmatrix}$, (12)
Where $c_{x}\equiv\cos(x/2)$ and $s_{x}\equiv\sin(x/2)$. This operator gives us the opportunity to generate transitions between all the basis states (in the computational basis) $\left\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\right\}$ depending on the $\theta$ and $\phi$ values.
For RFOX, XX remains at nearly unit strength throughout the circuit, while a harmonic ZX kick with zero time average acts as a first-order CD correction. Consequently, the schedule is not an adiabatic ramp of the form $(1-s)H_{\mathrm{driver}}+sH_{\mathrm{problem}}$; instead, it implements a fixed-depth quantum walk that exploits the wide XX non-stoquastic gap together with CD suppression of diabatic leakage.
For each edge $(u,v)\in E$ and each slice $k=0,\dots,p-1$, where $p$ denotes the number of time steps, we set
$\theta_{k}=1-\delta\cos\!\bigl(2\pi Nk/p\bigr),\quad\phi_{k}=\delta\sin\!\bigl(2\pi Nk/p\bigr),\quad k=0,\dots,p-1,$ (13)
with $\delta\ll 1$ (we fix $\delta=10^{-3}$) and $N$ the number of qubits. The choice of $\delta$ balances CD effects: it can be determined more precisely using the system’s energy gap, high-frequency convergence bounds, or a local sweep estimating the energy variance Sels and Polkovnikov (2017); Bukov et al. (2015). Numerical studies of XX catalysts in random Ising and Maximum-Weight Independent Set instances show typical gap amplifications of $2\times$ - $4\times$ compared to X-only drivers Feinstein et al. (2024). Because the diabatic runtime scales as $\Delta^{-2}$, this amplified gap alone offers a potential speed-up, even before CD corrections are applied.
III.3 Floquet-Magnus effective Hamiltonian
To derive the effective Hamiltonian, we apply the Floquet-Magnus expansion Blanes et al. (2009); Bukov et al. (2015); Kuwahara et al. (2016). Starting from an edge $(u,v)\in E$, define the Pauli bilinears
$H_{XX}=X_{u}X_{v},\quad H_{ZX}=Z_{u}X_{v}.$ (14)
The time‐dependent interaction Hamiltonian is:
$H_{\rm int}(t)=A_{XX}(t)\sum_{(u,v)\in E}H_{XX}+B_{ZX}(t)\sum_{(u,v)\in E}H_{ZX},$ (15)
with harmonic envelopes
$A_{XX}(t)=1-\delta\cos(\omega t),\quad B_{ZX}(t)=\delta\sin(\omega t)\ ,$ (16)
[…]
Adding the field-encoding Hamiltonian:
$H_{B}=\sum_{j=1}^{n}\phi_{j}X_{j}\ ,$ (17)
the full Hamiltonian reads
$H_{\rm full}(t)=H_{B}+\left(1-\delta\cos(\omega t)\right)H_{XX}+\delta\sin(\omega t)H_{ZX}\ .$ (18)
Using the Magnus expansion general formula we have that for a $T$-periodic Hamiltonian $H(t+T)=H(t)$, with $T=2\pi/\omega$, the one-period propagator $U_{per}\equiv\mathcal{T}\text{exp}\left[-i\int_{0}^{T}H(t)dt\right]$ can be written as $U_{per}=\text{exp}\left[-i\Omega\right]$ with
$\Omega=\Omega_{1}+\Omega_{2}+\mathcal{O}(\delta^{3})\ ,$ (19)
$\Omega_{1}=\int_{0}^{T}H(t)dt,\quad\Omega_{2}=\frac{1}{2}\int_{0}^{T}dt_{1}\int_{0}^{t_{1}}dt_{2}[H(t_{1}),H(t_{2})].$ (20)
Splitting off the oscillating part we have $H_{\rm osc}(t)=\delta\left[-\cos(\omega t)H_{XX}+\sin(\omega t)H_{ZX}\right]$ and the time-dependent Hamiltonian can be rewritten as
$H_{\rm full}(t)=H_{B}+H_{XX}+H_{\rm osc}(t)\ .$ (21)
For the first Magnus term we have that $\int_{0}^{T}\text{cos}(\omega t)dt=\int_{0}^{T}\text{sin}(\omega t)dt=0$, meaning that the oscillating part vanishes. The first Magnus term then becomes:
$\Omega_{1}=H_{B}+H_{XX}\ .$ (22)
For the second Magnus term we evaluate the commutator of $H_{XX}$ and $H_{ZX}$ given as:
$\left[H_{XX},H_{ZX}\right]=\sum_{(u,v)\in E}\left[X_{u}X_{v},Z_{u}X_{v}\right]=\sum_{(u,v)\in E}X_{u}X_{v}(Z_{u}X_{v})-Z_{u}X_{v}(X_{u}X_{v})=\sum_{(u,v)\in E}-2iY_{u}\ ,$ (23)
then we take the time-dependent parts
$\left[H_{osc}(t_{1}),H_{osc}(t_{2})\right]=\delta^{2}\sin\left[\omega(t_{1}-t_{2})\right]\left[H_{XX},H_{ZX}\right]=2i\delta^{2}\sin\left[\omega(t_{1}-t_{2})\right]Y_{u}\ .$ (24)
We omit the $\sum_{(u,v)\in E}$ part to simply the expression, but the term $Y_{u}$ maintains the summation over all edges. The second Magnus term is given by
$\Omega_{2}=\frac{1}{2}\int_{0}^{T}dt_{1}\int_{0}^{t_{1}}dt_{2}\ \left(-2i\delta^{2}\sin\bigl[\omega(t_{1}-t_{2})\bigr]\right)Y_{u}=-i\delta^{2}\Bigl[\frac{T}{\omega}\Bigr]\sum_{(u,v)\in E}Y_{u}.$ (25)
we have the effective Hamiltonian to $\mathcal{O}(\delta^{2})$:
$H_{\rm eff}=H_{B}+\sum_{(u,v)\in E}X_{u}X_{v}-i\frac{\delta^{2}}{\omega}\sum_{(u,v)\in E}Y_{u}+\mathcal{O}(\delta^{3}/w^{2})\ .$ (26)
RFOX maintains a robust non-stoquastic XX backbone that prevents gap closures, while a high-frequency ZX kick serves as a CD pulse Schindler and Bukov (2024); Claeys et al. (2019). In this way, the phase information encoding the RFIM cost accumulates without collapsing the spectral gap, and the final projective measurement produces a bit-string distribution sharply peaked at the ground state.
IV Problems to analyze
For simulations we generated multiple sets of 150 random graphs with associated magnetic fields on each node. Each set comprises two graph model generators: Erdös-Rényi and Watts-Strogatz, each graph set with $n\in\{7,9,12\}$ nodes. To each graph, we assigned node fields drawn uniformly from one of three ranges: [1,1][-1,1], [3,3][-3,3], or [5,5][-5,5], giving a total of 2700 RIM instances.
We also performed experiments on IBM Quantum hardware. On ibm_brisbane, we ran baseline analyses for our metrics. On ibm_sherbrooke and ibm_torino, we evaluated two sets of nine RFIM instances (one for each quantum computer), comprising three graph sizes and three field-range distributions, for each of the algorithmic models: RFOX, XX, and X+sXX; both in simulation and on hardware.
IV.1 Erdös-Rényi graphs
The first generator is called Erdös-Rényi (sub Fig. 1(a)) model $G(n,p)$ with $n\in\mathbb{N}$ number of vertices and vertex set $V=\{0,1,\dots,n-1\}$. Each unordered pair $\{u,v\}\subset V$, $u\neq v$, is included as an edge independently with probability $p\in[0,1]$ (for our simulations $p=0.8$). The probability of obtaining a specific graph $G=(V,E)$ is:
$\Pr\bigl(G_{n,p}=G\bigr)=p^{|E|}\,(1-p)^{\binom{n}{2}-|E|}\ .$ (27)
(a) RFIM using a 7-node Erdös-Rényi graph with $p=0.7$.
(b) RFIM using a 9-node Watts-Strogatz graph with $p=0.5$.
Figure 1: RFIM graph generators with magnetic field range [1,1][-1,1]. Red scales represent positive magnetic values and blue scales represent negative magnetic values.
IV.2 Watts-Strogatz graphs
The second generator is the Watts-Strogatz (sub Fig. 1(b)) model which can produce small-world graphs, which differ from the previous graph generator. Let $n\in\mathbb{N}$ be the number of vertices, $k\in\{2,4,6,\dots,n-1\}$ the even mean degree (we set $k=6$), and $p\in[0,1]$ the edge-rewiring probability. We begin with the regular ring lattice $R_{n,k}$ on vertex set $V=\{0,1,\dots,n-1\}$ and initial edge set
$E_{0}=\bigl\{\{i,(i+r)\bmod n\}\mid i\in V,\;r=1,\dots,\tfrac{k}{2}\bigr\}\ ,$ (28)
so that $|E_{0}|=\tfrac{nk}{2}$. Each clockwise edge $(i,r)$, with endpoint $j=(i+r)\bmod n$ and current edge set $E$, is rewired independently: with probability 1p1-p we keep $\{i,j\}$; and with probability $p$ we remove it and choose a new endpoint
$w\sim\mathrm{Unif}\bigl(V\setminus(\{i\}\cup N_{E}(i))\bigr),$ (29)
where $N_{E}(i)$ is the set of vertices adjacent to $i$. We then add the shortcut edge $\{i,w\}$. We set $p=0.7$ for our simulations, where the resulting graphs exhibit both high clustering and short average path lengths.
IV.3 Real quantum hardware problems
For the real-hardware experiments, we used the topologies (Figure 2) of IBM Quantum’s ibm_brisbane, ibm_sherbrooke and ibm_torino we tested RFIM sets employing between 12 and 20 physical qubits. We selected magnetic fields from a uniform distribution over the ranges [1,1][-1,1], [2,2][-2,2], or [3,3][-3,3].
Figure 2: RFIM instance on IBM Quantum hardware with 15 physical qubits and field values in [3,3][-3,3].
We limited the number of qubits to maintain feasibility on current NISQ hardware and to enable direct comparison with simulation results.
V Results Analysis
In this section, we compare the performance of RFOX against several baselines. For simulations, we benchmark against: (i) the X (stoquastic) conventional annealing schedule, used as a baseline; and (ii) two non-stoquastic drivers: an XX schedule and a combined X+sXX schedule.
Their annealing schedule is:
$H(s)=(1-s)H_{\rm driver}+s\,H_{\rm problem},\quad H_{\rm driver}=\begin{cases}\ H_{X}\\ H_{XX},\\ H_{X}+\ s\ H_{XX}\ ,\end{cases}$ (30)
with $J_{XX}=1$ for non-stoquastic drivers Choi (2021); Albash (2019), and the standard problem (cost) Hamiltonian given by $H_{\rm problem}=-\sum_{\langle i,j\rangle\in E}J_{ij}\,Z_{i}Z_{j}-\sum_{i}h_{i}\,Z_{i}$.
First, we analyze the spectral gap of two representative RFIM instances, one using a Erdös-Rényi graph and one a Watts-Strogatz graph, to illustrate how RFOX maintains its performance advantage over the X, XX and X+sXX drivers. Next, we present simulation results on Erdös-Rényi and Watts-Strogatz ensembles (see Section IV), computing for each instance:
  • Minimum cost per instance: the energy of the most frequently observed bitstring relative to the theoretical minimum;
  • Expected Energy Value (EEV): the ensemble‐averaged energy weighted by measurement frequencies;
  • Hamming distance per instance: the distance between the most frequent bitstring and the exact ground‐state bitstring.
We report both raw values and their tendency lines across each problem set, emphasizing minimum cost as the primary performance metric.
Finally, we describe hardware experiments on IBM Quantum backends (Eagle r3 and Heron r1), we compare RFOX against the two non-stoquastic approaches only. We apply analogous metrics, subject to runtime and noise constraints, to enable a direct comparison with the simulation results.
V.1 Energy gap analysis
In order to understand RFOX’s performance relative to conventional and non-stoquastic drivers, we analyze two example RFIM instances: an Erdös-Rényi graph with $N=7$ nodes and field range [3,3][-3,3], and a Watts–Strogatz graph with $N=9$ nodes and field range [2,2][-2,2] (see Figure 3).
(a) 6-node Erdös-Rényi [3,3][-3,3].
(b) 9-node Watts-Strogatz [2,2][-2,2].
Figure 3: RFIM instances used for energy-gap analysis.
For each Trotter slice $k=0,1,\dots,p-1$, we diagonalize the instantaneous Hamiltonian $H_{k}$:
$H_{k}|E_{l,k}\rangle=E_{l,k}|E_{l,k}\rangle,\quad l=0,1,\dots$ (31)
and define the gap
$\Delta_{k}=E_{1,k}-E_{0,k}\ ,\quad\Delta_{\min}=\min_{k}\Delta_{k}.$ (32)
The compared methods, expressed as slice-Hamiltonians for $k=0,1,\dots,p-1$, are:
$H_{k}^{\rm RFOX}=H_{B}+\theta_{k}\sum_{(u,v)\in E}X_{u}X_{v}+\phi_{k}\sum_{(u,v)\in E}Z_{u}X_{v},$ (33)
$H_{k}^{X}=(1-s_{k})\sum_{j}X_{j}+s_{k}\Bigl(\sum_{(u,v)\in E}Z_{u}Z_{v}+\sum_{j}h_{j}Z_{j}\Bigr),$ (34)
$H_{k}^{XX}=(1-s_{k})\sum_{(u,v)\in E}X_{u}X_{v}-s_{k}\Bigl(\sum_{(u,v)\in E}Z_{u}Z_{v}+\sum_{j}h_{j}Z_{j}\Bigr),$ (35)
$H_{k}^{X+sXX}=-(1-s_{k})\sum_{j}X_{j}+s_{k}(1-s_{k})\sum_{(u,v)\in E}X_{u}X_{v}-s_{k}\Bigl(\sum_{(u,v)\in E}Z_{u}Z_{v}+\sum_{j}h_{j}Z_{j}\Bigr)\ ,$ (36)
with $s_{k}=k/p\in[0,1]$ and $p=100$.
Figure 4: Energy gap for the 6-node Erdös-Rényi instance: RFOX, $X$-only, and X+sXX.
The gap analysis of the 7-qubit Erdös-Rényi instance (Fig. 4) reveals a clear hierarchy. RFOX maintains an almost flat gap of $\Delta_{k}\simeq 1.9-2.0$ across all slices. In contrast, the XX-only driver starts with an exponentially small gap ($\Delta_{\text{min}}\approx 0.03$) and recovers after the first third of the sweep; the X+sXX schedule dips to $\Delta_{\text{min}}\approx 0.17$ before climbing; and the conventional X-driver sees its gap collapse late to $\Delta_{\text{min}}\approx 0.18$. Because the adiabatic runtime scales as $\Delta^{-2}_{\text{min}}$, RFOX is expected to be roughly an order of magnitude faster than either X or X+sXX and three to four orders of magnitude faster than XX, highlighting the benefit of an almost constant non-stoquastic gap combined with counter-diabatic steering.
It is worth noting that, in this toy example, all four schedules perform reasonably well. The crucial difference, however, lies in gap stability. RFOX preserves a robust minimum gap even with an approximate choice of $\delta$. By contrast, the X-driver begins with a sizable gap but sees it steadily shrink as the evolution progresses; the XX driver starts with an almost vanishing gap, prolonging the period in which diabatic leakage is likely; and the X+sXX schedule suffers a sharp gap reduction in the first quarter of the evolution, mirroring the runtime penalty of the XX driver. These gap collapses translate into substantially longer evolution times.
Figure 5: Energy gap for the 9-node Watts–Strogatz instance: RFOX, XX‐only, and X+sXX.
For the 9-node Watts-Strogatz instance (Fig. 5), the gap profile again highlights RFOX’s robustness. RFOX maintains a nearly flat gap, oscillating gently around $\Delta_{k}\approx 2.5$ for the entire sweep. The X-only baseline starts with a moderate gap ($\approx 2.0$) but declines steadily to a late-time minimum of $\approx 0.5$, while the XX driver begins essentially gap-less and recovers only after $k/p\approx 0.35$, leaving a long, vulnerable interval. The X+sXX catalyst fares better than XX alone yet still plunges below $0.4$ early in the evolution before climbing above the RFOX curve near the end. Minimum-gap values are therefore $\Delta_{\min}^{\text{RFOX}}\approx 2.4$, versus $\approx 0.5$ for X, $\approx 0.05$ for XX, and $\approx 0.35$ for X+sXX. Given the $\Delta^{-2}_{\text{min}}$ scaling of the adiabatic runtime, RFOX is projected to be roughly an order of magnitude faster than the best stoquastic baseline and many orders faster than the $XX$ driver. Equally important, its flat profile eliminates the single bottleneck slice that dominates runtime in the other schedules, confirming that RFOX remains effective even on small-world topologies.
V.2 Results for Erdös-Rényi RFIM instances
For the Erdös-Rényi instances we fix $\delta=0.001$ and the number of time slices $p=100$ for every RFIM problem when running RFOX (blue). The same settings are applied to the other drivers: X (orange), representing the conventional stoquastic adiabatic approach; XX (green); and X+sXX (light blue), representing the non-stoquastic drivers.
For the Erdös-Rényi RFIM problems (Fig. 6), the results show that RFOX consistently outperforms the other drivers (X-stoquastic, XX, and X+sXX non-stoquastic) across multiple performance metrics. In terms of cost difference (sub Fig. 6(a), closer to zero is better) and minimum cost per instance (sub Fig. 6(d), lower better), RFOX not only achieves the smallest deviations from the optimal solutions but also maintains stable performance trends across different magnetic field ranges and graph sizes, it is important to mention that RFOX is able to achieve the real ground state of the problem more consistently compared to the other approaches. Notably, RFOX performs better in more complex problem instances, i.e., those with larger magnetic field ranges. This robustness suggests that the oscillatory exchange mechanism effectively adapts to varying problem scales, mitigating the degradation in solution quality often observed in other models. The EEV results (sub Fig. 6(b), lower is better) show a clear trend. Even in cases where RFOX does not achieve the lowest EEV on the smallest instances, its EEV steadily decreases as the number of qubits grows, indicating that RFOX gains relative advantage on more complex problems. In a few easy instances the stoquastic X driver is competitive, these are typically cases where near-trivial configurations (e.g., all spins aligned) maximize the interaction term. However, as the magnetic-field range increases and the instances become less trivial, RFOX consistently outperforms both the X driver and the non-stoquastic baselines.
The Hamming distance analysis (sub Fig. 6(c), closer to zero is better) reinforces these findings: RFOX consistently maintains small distances from the optimal bitstring, while other models, particularly X and X+sXX, exhibit broader distributions and larger deviations as the number of nodes and magnetic field range increase. This again arises from the fact that these models often converge to configurations dominated by interaction energy, resulting in trivial solutions where all spins (qubits) are aligned, e.g., $|00\cdots 0\rangle$ or $|11\cdots 1\rangle$, which can be the opposite of the true ground state. RFOX, by contrast, is more effective at preserving the correct solution structure during evolution, avoiding high-cost plateaus and large configuration jumps. Overall, these results demonstrate that RFOX not only excels in solution quality but also offers scalability and robustness advantages for highly dense graphs such as the Erdös-Rényi RFIM instances.
(a) Cost difference best vs. theoretical minimum.
(b) EEV per instance.
(c) Hamming distance best vs. theoretical minimum.
(d) Cost for most occurred state.
Figure 6: Results for 150 random instances of RFIM problems using Erdös-Rényi graphs with magnetic fields [1,1][-1,1], [3,3][-3,3] and [5,5][-5,5] with 7-, 9-, and 12-nodes. The x-axis shows the magnetic range for each of the generated RIM sets depending on the number of nodes.
V.3 Results for Watts-Strogatz RFIM instances
For the Watts–Strogatz RFIM problems (Fig. 7), the results show a consistent pattern where RFOX (blue) outperforms the other drivers across most metrics, though the margin of superiority depends on the problem complexity.
In terms of cost difference (sub Fig. 7(a)) and minimum cost per instance (sub Fig. 7(d)), RFOX achieves the lowest deviations from the theoretical optimum, with flatter tendency lines compared to the steep increases observed for the XX and X+sXX non-stoquastic drivers, and the X stoquastic driver. This stability across different magnetic field ranges and graph sizes suggests that RFOX is more resilient to structural irregularities inherent to the Watts-Strogatz topology, which often features clustered connectivity and shortcut edges.
The EEV results (sub Fig. 7(b)) reveal that for low magnetic field ranges, both RFOX and the XX-based drivers yield similar energy performance; however, as the range increases, the non-stoquastic and stoquastic drivers show more pronounced deviations. The Hamming distance (sub Fig. 7(c)) analysis further confirms RFOX’s robustness: it consistently yields solutions with smaller distances from the optimal bitstring, while the X and X+sXX drivers show a clear trend of increasing Hamming distances as both the number of nodes and field range grow. This reflects a tendency of the latter models to converge to spin configurations dominated by interaction alignment (e.g., all spins up or down), which can diverge significantly from the true ground state in these problems.
In general, the most informative performance indicators are the minimum cost per instance and the cost difference, as these directly measure the effectiveness of an optimization algorithm. For combinatorial optimization problems such as the RFIM, the primary objective is to produce a concrete optimal, or near-optimal, solution. In the case of RFOX, the algorithm can, in certain instances, reach the theoretical minimum and, in others, deliver a close approximation. These outcomes are particularly desirable because the goal is for the optimal solution to emerge as the most frequently observed state at the end of the quantum algorithm, similar in spirit to Grover’s algorithm, without requiring additional classical post-processing. By contrast, relying on the EEV is less practical for larger instances, as the number of measured states can grow substantially, a limitation that becomes even more pronounced on actual NISQ devices. Overall, these results demonstrate that RFOX not only scales effectively to the Erdös-Rényi and Watts-Strogatz problems structures but also maintains solution quality under increasingly complex conditions, offering a more balanced and adaptable optimization approach compared to the other tested drivers.
(a) Cost difference best vs. theoretical minimum.
(b) EEV per instance.
(c) Hamming distance best vs. theoretical minimum.
(d) Cost for most occurred state.
Figure 7: Results for 150 random instances of RFIM problems using Watts-Strogatz graphs with magnetic fields [1,1][-1,1], [3,3][-3,3] and [5,5][-5,5] with 7-, 9-, and 12-nodes. The x-axis shows the magnetic range for each of the generated RIM sets depending on the number of nodes.
V.4 Real hardware experiments on IBM Quantum
For the hardware experiments we adopted a common parameter set to keep the comparison with the simulations meaningful while respecting real-device constraints. We fixed the kick amplitude at $\delta=0.001$, matching the simulation setting, and limited every schedule to $p=50$ time slices. Although this depth is much shallower than the $p=100$ used in simulation, it tries to reduce cumulative gate errors and fits within IBM’s execution time limits.
We first apply the proposed metrics to a single 12-qubit RFIM instance on ibm_brisbane (Figure 8).
Figure 8: 12-qubit RFIM instance on ibm_brisbane using field values in [3,3][-3,3].
V.4.1 String-overlap fidelity
To quantify performance, we compute the string-overlap fidelity between the most frequently observed bitstring (the winner) and the theoretical ground-state bitstring (gs):
$F_{\mathrm{overlap}}=1-\frac{d_{H}(\mathrm{winner},\mathrm{gs})}{n}\ ,$ (37)
where $n=12$ and the Hamming distance given by
$d_{H}(g,w)=\sum_{i=1}^{n}(1-\delta_{g_{i},w_{i}})=\sum_{i=1}^{n}\lvert g_{i}-w_{i}\rvert$, (38)
with $g_{i},w_{i}\in\{0,1\}$. A fidelity $F_{\mathrm{overlap}}=1$ indicates that the winner exactly matches the ground state; lower values quantify the fraction of mismatched bits.
Figure 9: String-overlap fidelity (higher better) for the 12-qubit RFIM instance in simulation vs. hardware.
In the noiseless simulator RFOX reaches perfect overlap ($F_{\text{overlap}}=1$), while the slow-ramp X+sXX attains 0.74 and the stoquastic XX driver only 0.66. Hardware noise reduces all scores, but the ranking survives: RFOX-real still matches 58 % of the ground-state bits, outperforming XX-real (49 %) and X+sXX-real (33 %). The larger simulator–hardware drop for RFOX shows the current noise floor, yet even in this regime RFOX’s coherent, gap-protected evolution delivers the closest candidate state to the optimum, validating its practical advantage over both baseline drivers.
V.4.2 Distribution overlap via Jensen–Shannon distance
Continuing our analysis, we compute the Jensen–Shannon distance between the hardware output distribution $p$ and the simulated distribution $q$:
$D_{\mathrm{JS}}(p\parallel q)=\tfrac{1}{2}\bigl[D_{\mathrm{KL}}(p\parallel m)+D_{\mathrm{KL}}(q\parallel m)\bigr]\ ,\quad m=\tfrac{1}{2}(p+q)\ ,$ (39)
where the Kullback-Leibler divergence is $D_{\mathrm{KL}}(k\parallel m)=\sum_{x\in\mathcal{X}}k(x)\,\log\!\frac{k(x)}{m(x)}$, with $k$ equal to $p$ or $q$.
Figure 10: Jensen–Shannon distance (lower better) for the 12-qubit RFIM instance in simulation vs. hardware.
The Jensen–Shannon distance quantifies how much the hardware output deviates from the ideal, noiseless simulation. In our 12-qubit experiment, $D_{\mathrm{JS}}=0.888$ for RFOX, whereas the $XX$ and slow‐ramp X+sXXX+s\,$X$ protocols yield $0.985$ and $0.994$, respectively (Fig. 10). Since $D_{\mathrm{JS}}=0$ indicates perfect agreement and $D_{\mathrm{JS}}=1$ a complete statistical mismatch, these values show that RFOX retains substantially more of the ideal distribution under real‐device noise than either baseline.
V.4.3 Average Hamming distance to the optimum
To quantify how far the output distribution lies from the true ground state $x_{\min}$, we compute the average Hamming distance:
$\langle d_{H}\rangle=\sum_{x}p_{x}\,d_{H}(x,x_{\min}),$ (40)
where $p_{x}$ is the probability of measuring bitstring $x$ and $d_{H}(x,x_{\min})$ is the Hamming distance to the optimal bitstring. Figure 11 displays the resulting distributions.
Figure 11: Average Hamming distance (lower better) to the optimum for the 12-qubit RFIM instance in simulation vs. hardware.
In ideal simulation, RFOX finishes closest to the optimum (lower better) with an average of 2.1 bit flips, while the slow-ramp X + sXX and stoquastic XX drivers average 3.2 and 3.5 flips, respectively. On IBM Quantum hardware, noise inflates all distances, yet RFOX still leads with 5.8 flips compared to 6.3 for both baselines.
V.4.4 RFIM experiments on ibm_sherbrooke and ibm_torino
For the final results, we generated random RFIM instances on the IBM Quantum backends ibm_sherbrooke and ibm_torino, using $N\in\{12,15,20\}$ physical qubits and field ranges [1,1][-1,1], [2,2][-2,2], and [3,3][-3,3]. We fixed the CD amplitude at $\delta=0.001$ (matching the simulation) and employed $p=50$ time-slices for both simulated and hardware runs. This expanded dataset provides a more comprehensive evaluation of each algorithm’s performance under realistic noise conditions.
(a) ibm_sherbrooke
(b) ibm_torino
Figure 12: String-overlap fidelity (higher better) for hardware vs. simulation using ibm_sherbrooke and ibm_torino quantum computers.
For the string-overlap fidelity results showed in Figure 12 both hardware runs reproduce the qualitative ranking observed in simulations, but the absolute fidelities now depend strongly on the underlying device family. On ibm_sherbrooke (Eagle r3) the winner-string overlap of RFOX stabilizes around overlap $F_{\text{overlap}}\simeq 0.5-0.55$ for every instance, remaining essentially flat as disorder and problem size increase; the XX driver stays 5-10 points lower for the larger-field tasks, while X+sXX hovers in between. On ibm_torino system (Heron r1, lower-noise silicon spin qubits) lifts all hardware bars by roughly 10% points, with RFOX reaching 0.60-0.65 and the two baselines closing, but not eliminating, the gap to the simulator. Crucially, RFOX maintains the smallest hardware-measured separation from the theoretical ground state, showing that RFOX can exploit windows of improved coherence and sustain robustness under elevated noise. On the Heron r1 processor, this translates into a substantial increase in the fidelity of the measured state to the true ground-state.
(a) ibm_sherbrooke
(b) ibm_torino
Figure 13: Jensen–Shannon distance (lower better) for hardware vs. simulation using ibm_sherbrooke and ibm_torino quantum computers.
The Jensen-Shannon analysis (Figure 13) confirms that RFOX preserves the simulator distribution substantially better than the two baselines on both devices. On ibm_sherbrooke the distance lies in the range $D_{JS}=0.82-0.97$, whereas XX and X+sXX schedules are almost fully decorrelated from their noiseless counterparts ($D_{JS}\gtrsim 0.97$ for every hard instance). Then, ibm_torino shifts all bars downward by $\approx 0.05$, yet RFOX remains the best‑overlapping path, now at $D_{JS}=0.78-0.92$ versus 0.930.990.93-0.99 for the baselines. Thus the RFOX shallow-depth design not only tolerates the higher error rates of the Eagle architecture but also continues to deliver the closest hardware-simulator agreement when coherence improves, underscoring the intrinsic noise resilience of the RFOX schedule.
(a) ibm_sherbrooke
(b) ibm_torino
Figure 14: Average Hamming distance (lower better) to the optimum for hardware vs. simulation using ibm_sherbrooke and ibm_torino quantum computers.
The average Hamming distances results presented in the Figure 14 we see that on ibm_sherbrooke RFOX achieves the smallest hardware distance in every problem class, typically $\langle d_{H}\rangle=6-7$ for the 12- and 15-qubit sets and $\langle d_{H}\rangle\approx 10$ for 20 qubits, while the XX-only and X+sXX schedules overshoot by 20-30%. The simulator bars (pale tones) sit 1-2 flips lower than hardware for RFOX but 3-4 flips lower for the baselines, showing that the constant-gap walk converts coherence into solution quality more efficiently. ibm_torino improves all hardware distances by roughly two bit flips; RFOX now lands at $\langle d_{H}\rangle=5-6$ (12-15 qubits) and $\approx 9.5$ (20 qubits), keeping its lead over both baselines, which remain 15–25% farther from the optimum. RFOX present a better reduction in the hardware-simulator gap than the other methods, confirming that the schedule’s harvest the additional coherence of the Heron architecture more effectively.
These hardware results underscore two key observations. First, on low-density RFIM graphs, conventional non-stoquastic drivers XX and X+sXX) perform better (for simulation cases) than on the denser graphs used in simulation, yet they still fall short of RFOX’s consistency. Second, while the Eagle r3 processor exhibits significant cumulative errors that constrain circuit depth, experiments on the Heron r1 backend deliver markedly improved solution quality, especially for RFOX, consistent with our simulation results. Altogether, RFOX maintains a decisive performance advantage in both simulation and on real quantum hardware.
In both the baseline run on ibm_brisbane and the subsequent experiments on ibm_sherbrooke, which have the same type of quantum processor, RFOX completes each RFIM instance in roughly 7 seconds, whereas the XX and X+sXX schedules require about 20 seconds. On ibm_torio RFOX experiments took around 4-5 seconds where the other methods took around 7-9 seconds, we see an improvement in the execution time in all models where RFOX maintain the lowest execution time of all. This threefold speed-up confirms that reformulating the RFIM via magnetic-field encoding—combined with an almost constant non-stoquastic term and counter-diabatic steering—yields not only higher solution accuracy but also noticeably shallower, faster-running circuits.
VI Conclusions
Our numerical study demonstrates that the RFOX schedule—an always-on, non-stoquastic XX catalyst augmented by a weak, harmonic ZX counter-diabatic kick—consistently outperforms both the X-only and XX drivers, and the commonly used X+sXX catalyst. Across all tested random-field Ising model instances (7, 9, and 12 qubits; three magnetic-field ranges; Erdös–Rényi and Watts–Strogatz graphs), RFOX achieves the lowest cost average, the smallest excess energy above the exact optimum, and the shortest average Hamming distance to the ground-state bitstring. These performance gains directly reflect the Floquet-Magnus analysis: the first-order term preserves the full strength of the non-stoquastic XX driver, providing a robust gap floor, while the second-order term introduces only a perturbative $Y$-field micromotion counter-diabatic correction that refines, rather than narrows, the spectrum. Consequently, the minimum gap remains nearly constant throughout the evolution, whereas comparison schedules exhibit slow starts, gap shrinkage, or even pronounced gap collapse. Since adiabatic runtime scales as $\Delta_{\min}^{-2}$, this nearly flat-gap profile enables RFOX to approach the ground state in far fewer Trotter slices, and thus in shallower circuits, than its competitors.
Hardware experiments on Eagle-class (ibm_brisbane and ibm_sherbrooke) and Heron-class (ibm_torino) backends confirm that RFOX’s advantage persists under realistic noise. Although decoherence increases errors across all metrics, RFOX consistently outperforms baseline schedules. This robustness is partly attributable to amplitude-damping noise, which shifts probability weight toward low-energy configurations, a shift that RFOX’s large minimum gap helps preserve. Importantly, RFOX delivers these improvements with shallower circuits as those of the comparator methods and requires no classical parameter optimization.
Two further observations arise from our hardware tests. First, on low-density RFIM graphs, conventional non-stoquastic drivers (XX and X+sXX) perform better than they do on the denser graphs studied in simulations. Second, current NISQ devices (e.g., IBM Eagle r3 and Heron r1) accumulate significant errors that severely limit feasible circuit depth. Nonetheless, RFOX’s consistent performance advantage highlights its potential as a scalable, noise-resilient quantum optimizer.
Overall, our findings indicate that a near constant-gap, non-stoquastic driver enhanced by a perturbative counter-diabatic term offers an effective and hardware-friendly strategy for quantum optimization. Future research should explore adaptive schemes for the kick amplitude, extend RFOX to a broader class of combinatorial problems, and integrate it with advanced error-mitigation techniques to push beyond the current $\sim$20-qubit scale.
VII Future work
The present study establishes RFOX as a promising, hardware-friendly schedule for random-field Ising models, yet several research directions remain open.
  • Adaptive counter-diabatic control.
  • Alternative non-stoquastic drivers.
  • Broader problem classes.
  • Analytic complexity bounds.
  • Qudit and higher-dimensional generalizations.
Pursuing these directions will clarify the ultimate scalability of constant-gap, non-stoquastic schedules and help turn RFOX from a proof-of-principle into a practical tool for quantum optimization in the NISQ era and beyond.
VIII Data availability
The code, raw graphs and histograms can be consulted in the following link: https://github.com/BrianSarmina/Papers/tree/main/RFOX .
IX Acknowledgments
Brian García Sarmina thanks Prof. T. Albash for valuable discussions and clarifications on non-stoquastic Hamiltonians and their role in adiabatic quantum evolution within the RFOX framework.
Appendix A Magnetic field phase mapping
We start from the all-zero state $|\psi_{0}\rangle=\bigotimes_{i=1}^{n}|0\rangle$ and create a uniform superposition over all computational basis states by applying the Walsh-Hadamard transform to each qubit:
$|\psi_{1}\rangle=H^{\otimes n}|\psi_{0}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{x=0}^{2^{n}-1}|x\rangle.$ (41)
Here, $|x\rangle$ runs over all bit-strings of length $n$, ensuring equal amplitude sampling of every spin configuration.
To encode local fields $h_{i}$, we apply a phase gate on each qubit:
$P(\phi_{i})=\begin{pmatrix}1&0\\ 0&e^{i\phi_{i}}\end{pmatrix}$, (42)
which imprints a relative phase $e^{i\phi_{i}}$ on the $|1\rangle$ component. We choose $\phi_{i}\propto h_{i}$ by normalizing $h_{i}\in[-h_{\max},h_{\max}]$ into $0,\pi]$ via
$\phi_{i}=\pi\,\frac{h_{i}/|h|_{\max}+1}{2}.$ (43)
After the phase embedding, the state becomes:
$|\psi_{2}\rangle=\bigotimes_{i=1}^{n}P(\phi_{i})\,|\psi_{1}\rangle=\frac{1}{\sqrt{2^{n}}}\sum^{2^{n}-1}_{x=0}e^{i\sum_{k=1}^{m}b_{k}\phi_{k}}\,|x\rangle\ ,$ (44)
where each basis state $|x_{i}\rangle=|b_{1}\dots b_{m}\rangle$ acquires the phase $\sum_{k}b_{k}\phi_{k}=\Phi_{x}$.
A.1 Phase encoding interference
After encoding the local fields with phase gates, we apply a second layer of Hadamard gates $H^{\otimes n}$. This step converts the encoded phases into amplitude variations, creating interference patterns that highlight the field contributions in the measurement basis. Concretely, if the system is initially in the state:
$|\psi_{2}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{x=0}^{2^{n}-1}e^{i\Phi_{x}}\,|x\rangle\ ,$ (45)
where each $\Phi_{x}$ depends on $\{\phi_{i}\}$, then:
$|\psi_{3}\rangle=H^{\otimes n}|\psi_{2}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{x=0}^{2^{n}-1}e^{i\Phi_{x}}\bigl(H^{\otimes n}|x\rangle\bigr)\ .$ (46)
Using
$H^{\otimes n}|x\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{y=0}^{2^{n}-1}(-1)^{x\cdot y}|y\rangle,$ (47)
where $x\cdot y$ is the bitwise dot product (mod 2), we obtain
$H^{\otimes n}|\psi_{3}\rangle=\frac{1}{2^{n}}\sum_{x=0}^{2^{n}-1}e^{i\Phi_{x}}\left[\sum_{y=0}^{2^{n}-1}\left(-1\right)^{x\cdot y}|y\rangle\right]\ .$ (48)
Each basis state $|y\rangle$ thus acquires contributions from all $|x\rangle$, weighted by $e^{i\Phi_{x}}$ and $(-1)^{x\cdot y}$, producing constructive or destructive interference patterns determined by the phase distribution.
Translating this encoding into an effective Hamiltonian, we note
$H^{\otimes n}P(\phi_{j})H^{\otimes n}=e^{-i\frac{\phi_{j}}{2}HZH}=e^{-i\frac{\phi_{j}}{2}X},$ (49)
since $P(\phi)=e^{i\phi/2}e^{-i(\phi/2)Z}$ and the global phase $e^{i\phi/2}$ can be dropped. The resulting field Hamiltonian acts along the XX‐axis:
$H_{B}=\sum_{j=0}^{N-1}\phi_{j}X_{j},$ (50)
where $\phi_{j}=\pi\,(h_{j}/|h|_{\max}+1)/2$ encodes the raw field values $h_{j}$.